# NeuralRemaster: Phase-Preserving Diffusion for Structure-Aligned Generation

### v2

@@ A.1 Training and Inference Details
We start from the officially released checkpoints of SD 1.5 [29], FLUX-dev [18], and Wan2.2-14B [33], and finetune each model with phase-preserving noise. For SD 1.5, we experiment with both full finetuning and LoRA finetuning, while for FLUX-dev and Wan2.2-14B we use LoRA finetuning due to computational constraints. At inference time, for Wan2.2-14B we adopt the 4-step LoRA from LightX2V [^3] https://huggingface.co/lightx2v/Wan2.2-Lightning and apply it directly on top of our finetuned LoRA weights to accelerate sampling.
@@
where $\lambda>0$ is the rate parameter of the exponential distribution and $r_{0}>0$ controls the minimum cutoff. In our experiments, we set $\lambda=0.1$ empirically. We set the transition bandwidth parameter $\sigma=2$, which controls the smoothness of the frequency mask $M(u,v)$ around the cutoff radius $r$.
Figure 10: Visual realism measured by Appearance Score.
Figure 11: Structural alignment measured by depth SSIM.
Appendix B Ablation Studies
We ablate the choices of $r_{0}$, the minimal cutoff radius at training time, and the inference-time cutoff radius $r$ in this section. All ablation experiments are conducted using SD 1.5 with LoRA finetuning and evaluated on $1,000$ randomly selected samples from the UnrealCV test set. Figure 11 shows the Appearance Score (Sec. 4.1.3), which measures the photorealism of re-rendered images. Figure 11 shows the depth consistency measured by SSIM, reflecting how well the re-rendered image aligns structurally with the original input.
From both figures, we observe that increasing the cutoff radius $r$ at inference time improves structural alignment at the cost of reduced photorealism. The minimal cutoff threshold $r_{0}$ during training also affects performance across different inference-time radii $r$. A higher $r_{0}$ during training leads to better performance with higher $r$ during inference, making it suitable for re-rendering images of decent quality where the model primarily refines details. Conversely, a lower $r_{0}$ favors scenarios with smaller inference-time $r$, performing better on low-quality inputs that require larger visual changes to achieve photorealism. In our experiments, we set $r_{0}=4$ for a balanced trade-off between structure preservation and photorealism.

### v3

@@ 4.1.3 Training and Inference Details
We start from the officially released checkpoints of SD 1.5 [Rombach_2022_CVPR], FLUX-dev [flux2024], and Wan2.2-14B [wan2025], and finetune each model with phase-preserving noise. For SD 1.5, we experiment with both full finetuning and LoRA finetuning, while for FLUX-dev and Wan2.2-14B we use LoRA finetuning due to computational constraints. At inference time, for Wan2.2-14B we adopt the 4-step LoRA from LightX2V [^2] https://huggingface.co/lightx2v/Wan2.2-Lightning and apply it directly on top of our finetuned LoRA weights to accelerate sampling. FLUX and SD 1.5-based models are trained on PhotoConceptBucket dataset [^3] https://huggingface.co/datasets/bghira/photo-concept-bucket; WAN2.2-14B based models are trained on OpenSoraPlan dataset [^4] https://huggingface.co/datasets/LanguageBind/Open-Sora-Plan-v1.1.0/tree/main.
@@
where $\lambda>0$ is the rate parameter of the exponential distribution and $r_{0}$ controls the minimum cutoff. In our experiments, we set $\lambda=0.1$ empirically. We set the transition bandwidth parameter $\sigma=2$, which controls the smoothness of the frequency mask $M(u,v)$ around the cutoff radius $r$.
4.2 Results
Table 1: Quantitative evaluation results for photorealistic re-rendering on UnrealCV.
Prompt Align. Structure Alignment
Method CLIP $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ ABSREL $\downarrow$
ControlNet-Tile 0.3018 0.5831 0.6804 0.9484
PNP 0.2978 0.4076 0.6993 0.8870
FBSDiff 0.3255 0.3558 0.6792 0.9362
SDEdit 0.2989 0.4540 0.6754 0.8932
Ours 0.3212 0.2397 0.7481 0.5589
Table 2: Quantitative evaluation for stylized re-rendering.
Prompt Align. Structure Alignment
Method CLIP $\uparrow$ LPIPS $\downarrow$ SSIM $\uparrow$ ABSREL $\downarrow$
ControlNet-Tile 0.3022 0.5849 0.7508 0.9085
PNP 0.2983 0.3029 0.7796 0.8816
FBSDiff 0.3091 0.3150 0.7562 0.9736
SDEdit 0.3011 0.3456 0.7588 0.9148
Ours 0.3017 0.2759 0.7842 0.8087
Table 3: Efficiency comparison in extra parameters and FLOPs relative to the base model as well as wall-clock time.
ControlNet SDEdit PNP FBSDiff Ours
Extra Params +50% +0% +0% +0% +0%
Extra FLOPs +50% +0% +100% +1100% +0%
Time (s/image) 28.02 20.31 20.40 133.8 20.22
Photorealistic Re-Rendering. Quantitative results on UnrealCV re-rendering are summarized in Table 1, where all methods are implemented using the SD 1.5-based models for fair quantitative comparison. We evaluate SDEdit with Gaussian noise (SDEdit) and structured noise (Ours) at identical noise levels (t=0.5t{=}0.5). This comparison shows that, compared to Gaussian noise, structured noise consistently improves structure preservation (near 90% improvements on LPIPS) while maintaining text prompt alignment. Qualitative examples are shown in Figure 6, which compares our Flux-based model against stronger recent models such as FLUX-Kontext, and Qwen-Edit. Across both settings, all methods improve photorealism-as reflected by higher AS scores than the input images while $\phi$-PD achieves the highest photorealism and superior structure alignment. We observe that QWen-Edit produces visually high-quality results but often fails to maintain structural alignment with the input image; for example, in the first four cases, it enlarges the main subjects significantly. FLUX-Kontext aligns better with the input structure but provides only limited improvement in visual quality. Our method achieves both high visual fidelity and consistent structural alignment across frames.
Stylized Re-rendering. Results of stylized re-rendering are presented in Table 2, with representative examples shown in Figure 7. All models are based on SD 1.5. This task evaluates the model’s ability to alter appearance while preserving scene structure. As shown, $\phi$-PD produces visually coherent stylizations that maintain object boundaries and spatial consistency, while prior methods often distort geometry or introduce texture misalignment. Quantitatively, $\phi$-PD achieves similar prompt alignment and significantly higher structure alignment.
Input
Refer to caption Refer to caption Refer to caption
Cosmos-Transfer 2.5
Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption
Figure 9: Video re-rendering results on CARLA.
Efficiency Comparison. Table 3 compares the extra parameters and FLOPs relative to the base model, as well as the wall-clock time each method takes to process an image. ControlNet Tile adds 50% parameters and computation overhead (UNet + ControlNet model per step. PNP is 2 $\times$ more expensive than the base model due to inversion and sampling, however achieves similar wall-clock time thanks to optimized implementation from diffuser. FBSDiff is extremely expensive (12 $\times$ base model FLOPs) due to 1000 inversion steps and 100 sampling steps. Our method does not introduce extra parameters or FLOPs.
Figure 10: LPIPS and CLIP scores with different $r_{0}$ and $r$.
5 Ablation Studies
5.1 Cutoff Radius
We ablate the choices of $r_{0}$, the minimal cutoff radius at training time, and the inference-time cutoff radius $r$ in this section. All ablation experiments are conducted using SD 1.5 with LoRA finetuning and evaluated on $1,000$ randomly selected samples from the UnrealCV test set. Figure 10 shows how $r_{0}$ and $r$ affects these metrics. Increasing the inference-time cutoff radius $r$ significantly improves structural alignment: LPIPS drops from 0.440.450.44-0.45 to 0.300.380.30-0.38, while the change in text alignment is small: CLIP scores stay around 0.300.320.30-0.32. The minimal cutoff threshold $r_{0}$ during training influences the structural alignment with a large inference time cutoff radius $r$. It also affects performance across different inference-time radii $r$. A higher $r_{0}$ during training leads to better performance with higher $r$ during inference, while a lower $r_{0}$ favors scenarios with smaller inference-time $r$.
5.2 Phase Preservation in Latent Space.
Classical results on phase-structure correspondence [oppenheim1981importance] were established in pixel space, while modern diffusion models operate in VAE latent space. We apply phase preservation directly to latents. While a complete theoretical analysis of how phase-structure relationships transfer through nonlinear encoders is beyond the scope of this work, we provide empirical validation in Figure 11.
Figure 11: Phase-structure correspondence in VAE latent space.
We encode two images $I_{1}$ and $I_{2}$ into VAE latent representations $z_{1}$ and $z_{2}$, then swap their Fourier magnitude and phase components. Combining the phase from $z_{1}$ with the magnitude from $z_{2}$ produces a decoded image that preserves the structure of $I_{1}$ (the dog’s silhouette), as shown in the top row. The bottom row shows the reverse combination which preserves the structure of $I_{2}$ (road, horizon, vehicles). This suggests that VAE latents, which are trained to preserve perceptually meaningful spatial structure, maintain sufficient phase-structure correspondence for our method to be effective.
6 Conclusion
We introduced Phase-Preserving Diffusion ($\phi$-PD), a simple yet effective reformulation of the diffusion process that replaces Gaussian noise with structured noise, preserving image phase while randomizing the magnitude in the frequency domain. This simple change retains spatial alignment throughout sampling without modifying the architecture, altering training objectives, or introducing inference-time overhead. We also introduced Frequency-Selective Structured (FSS) noise, which provides continuous control over structural alignment rigidity through a single frequency cutoff parameter, making it broadly applicable to different applications.